# Text Animation Illustration

### v2

@@ 5.2. Ablation Study
We conducted an ablation study to analyze the contribution of each component in our proposed method: learnable base shape, legibility regularization, mesh-based structure preservation regularization, and frequency encoding with annealing. Visual results in Fig. 9 showcase the qualitative impact of removing each component. Quantitative results in Tab. 2 further confirm their effectiveness.
In addition to Perceptual Input Conformity (PIC) and Text-to-Video Alignment (X-CLIP score), we employed warping error to assess temporal consistency, following EvalCrafter (Liu et al., 2023). This metric estimates the optical flow between consecutive frames using the pre-trained RAFT model (Teed and Deng, 2020) and calculates the pixel-wise difference between the warped image and the target image. The lower warping error indicates smoother and more temporally consistent animations.
Base Shape The calculation of legibility and structure preservation regularization involves the base shape. Hence, when removing the learnable base shape, the legibility loss $\mathcal{L}_{legibility}$ is computed between every output frame and the input letter, while the structure preservation loss $\mathcal{L}_{structure}$ is applied between every pair of consecutive frames.
As observed in Fig. 9 (row 2), removing the shared learnable base shape results in inconsistent animations. Specifically, as highlighted by the red circle, the appearance of the bullfighter deviates significantly between frames, harming legibility. The finding is also supported by Tab. 2 (row 2), where removing the base shape results in significant degradation under all three metrics.
Legibility Regularization Without the perceptual regularization on the base shape, the base shape struggles to preserve legibility. As a result, each animation frame loses the letter “R” shape in Fig. 9 (row 3), leading to lower PIC in Tab. 2 (row 3).
Structure Preservation Regularization Removing mesh-based structure preservation allows the base shape’s structure to deviate from the original letter, causing the discontinuity between the bullfighter and cape in the base shape and all frames, as highlighted in Tab. 2 (row 4). Without this regularization term, the animation shows inconsistent appearances across different frames, which degrades the legibility, leading to the lowest PIC in Tab. 2 (row 4).
Frequency Encoding and Annealing When removing frequency encoding and coarse-to-fine annealing, the motion and geometry quality suffers. For example, the bullfighter animation in Fig. 9 (row 5) shows unreasonable motion and geometry details, resulting in an animation that does not accurately represent the text prompt. Moreover, the degradation in motion quality also harms the temporal consistency, Tab. 2 (row 5).
@@
Figure 9. Visual comparisons of ablation study. Removing base shape or structure preservation regularization results in shape deviation and flickering. Without legibility regularization, each animation frame loses the letter “R” shape. The absence of frequency encoding and annealing leads to the degradation of the motion quality and geometry details.
@@
{tblr}
cells=halign=c,valign=m, column1=halign=l, hline1,3,8=1-7, hline1,3,8=1.0pt, vline2,3,4=1-7, cell11=r=2,   Method & Optical Flow Perceptual Text-to-Video
[…]
Full Model ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.01645}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.5310}}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{21.4447}}$
@@
No Freq. $0.02222$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{0.5377}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{20.8280}}$
Table 2. Quantitative results of the ablation study. The best and second-best scores for each metric are highlighted in red and blue respectively.
5.3. Generalizability
@@
Our optimization framework, leveraging Score Distillation Sampling (SDS), achieves generalization across various diffusion-based text-to-video models. To demonstrate this, we applied different base models for computing $\mathcal{L}_{SDS}$, including the 1.7-billion parameter text-to-video model from ModelScope (Wang et al., 2023), AnimateDiff (Guo et al., 2024), and ZeroScope (Luo et al., 2023). Fig. 10 presents visual results for the same animation sample (“Knight”) with each base model.
While the letter “K” exhibits deformations and animation styles unique to each model, all animations accurately depict the user’s prompt and maintain the basic “K” shape. This showcases the generalizability of our method. Hence, future advancements in text-to-video models with stronger prior knowledge will benefit our approach.
@@
Figure 10. Visual results of the same animation sample using different text-to-video base models.
6. Conclusion
@@
We propose an automated text animation scheme, termed “Dynamic Typography,” that deforms letters to convey semantic meaning and animates them vividly based on user prompts. Our method is an end-to-end optimization-based approach and is generalizable to arbitrary words and motion patterns. Nevertheless, there remain several limitations. First, the motion quality can be bounded by the video foundation model, which may be unaware of specific motions in some cases. However, our framework is model-agnostic, which facilitates integration with future advancements in diffusion-based video foundation models. Besides, challenges arise when user-provided text prompts deviate significantly from original letter shapes, complicating the model’s ability to strike a balance between generating semantic-aware vivid motion and preserving the legibility of the original letter. We hope that our work can open the possibility for further research of semantic-aware text animation that incorporates the rapid development of video generation models.

### v3

@@ 5.2 Ablation Study
We conduct ablations to analyze the contribution of each component in our proposed method: learnable base shape, legibility regularization, and mesh-based structure preservation regularization. Visual results in Fig. 9 showcase the qualitative impact of removing each component. Quantitative results in Tab. 2 further confirm their effectiveness.
@@
Figure 9: Visual comparisons of ablation study. Some artifacts are highlighted in red cycles. Removing base shape or structure preservation regularization results in shape deviation and inconsistent artifacts. Without legibility regularization, each animation frame loses the letter “R” shape.
In addition to Perceptual Input Conformity (PIC) and X-CLIP score (Text-to-Video Alignment), we employ warping error to assess temporal consistency, following EvalCrafter [27]. This metric estimates the optical flow between consecutive frames using the RAFT model [50] and calculates the pixel-wise difference between the warped image and the target image. The lower warping error indicates smoother and more temporally consistent animations.
The calculation of legibility and structure preservation regularization loss involves the base shape. Hence, when removing the learnable base shape, the legibility loss $\mathcal{L}_{\text{legibility}}$ is computed between every output frame and the input letter, and the structure preservation loss $\mathcal{L}_{\text{structure}}$ is only applied between every pair of consecutive frames.
@@
Full Model $0.01645$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{0.5310}}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{21.4447}}$
@@
Table 2: Quantitative results of the ablation study. The best score for each metric is highlighted in red. The full model gets the best in PIC, text-to-video alignment, and second-best in optical flow warping error, indicating the effectiveness of each module.
Base Shape: As observed in Fig. 9 (row 2), removing the shared learnable base shape results in inconsistent animations. Specifically, as highlighted by the red circle, the appearance of the bullfighter deviates significantly between frames, harming legibility. The finding is also supported by Tab. 2 (row 2), where removing the base shape results in significant degradation under all three metrics.
Legibility Regularization: Without the perceptual regularization on the base shape, the base shape struggles to preserve legibility. As a result, each animation frame loses the letter “R” shape in Fig. 9 (row 3), leading to lower PIC in Tab. 2 (row 3).
Structure Preservation Regularization: Removing mesh-based structure preservation allows the base shape’s structure to deviate from the original letter, causing the discontinuity between the bullfighter and cape in the base shape and all frames, as highlighted in Fig. 9 (row 4). Without this regularization term, the animation shows inconsistent appearances across different frames, degrading legibility and leading to the lowest PIC in Tab. 2 (row 4).
5.3 Generalizability
@@
Our framework, leveraging Score Distillation Sampling (SDS), achieves generalization across various diffusion-based text-to-video models. To demonstrate this, we apply different base models for computing $\mathcal{L}_{\text{SDS}}$, including the 1.7-billion parameter text-to-video model from ModelScope [51], AnimateDiff [15], and ZeroScope [29]. Fig. 10 presents visual results for the same animation sample (“K” in “Knight”) with each base model.
@@
Figure 10: Visual results of the same animation sample (“K”) using different text-to-video base models. All generated animations accurately depict the user’s prompt and maintain the “K” shape, showing the generalizability of our methods across different text-to-video foundation models.
While the animation exhibits deformations and animation styles unique to each model, all animations accurately depict the user’s prompt and maintain the “K” shape. This showcases the generalizability of our method. Hence, future advancements in text-to-video models with stronger prior knowledge will benefit our approach.
6 Conclusion
@@
We propose a text animation scheme, termed “Dynamic Typography”, that deforms letters to convey semantic meaning and animates them vividly based on user prompts. To automate text animation generation, we propose an end-to-end optimization-based framework that leverages the video diffusion prior. Our method faithfully depicts the user’s prompt in animation while preserving the input letter’s legibility, and is generalizable to different fonts, prompts, and languages. Nevertheless, there remain several limitations. First, the motion quality can be bounded by the video foundation model, which may be unaware of specific motions in some cases. Luckily, our framework is model-agnostic, facilitating integration with future diffusion-based video foundation model advancements. Besides, challenges arise when user-provided text prompts deviate significantly from original letter shapes, complicating the model’s ability to strike a balance between generating semantic-aware vivid motion and preserving the legibility of the original letter. We hope that our work can open the possibility for further research of semantic-aware text animation that incorporates the rapid development of video generation models.
Appendix
Appendix A Implementation Details
The base field and motion field are jointly optimized. Following Gal et al. [14], we interleavely optimize local motion and global motion. Adam optimizer is adopted, and the learning rates are set to be $35e-$ 3, $35e-$ 3, and $31e-$ 3 to learn the base field, local motion, and global motion. For the global motion, we set the scaling factor to be $2.0$, $21e-$ 2, $25e-$ 2, and $11e-$ 1 for translation, rotation, scale, and shear.
We set the regularization weight to be $35e$ 3 for the $\mathcal{L}_{\text{legibility}}$ and $31e$ 3, $41e$ 4 for the $\lambda_{1}$, $\lambda_{2}$ in $\mathcal{L}_{\text{structure}}$. We observe that $\mathcal{L}_{\text{legibility}}$ often plays a dominant role. When perceptual regularization is employed from the start, the base shape typically retains the original letter’s form, preventing any semantic deformations. Hence, following Iluz et al. [19], we gradually increase the weight of $\mathcal{L}_{\text{legibility}}$ to make its effects after semantic deformation has taken place.
We use the text-to-video-ms-1.7b model in ModelScope [51, 29] for the diffusion backbone. We apply augmentations including random crop and random perspective to all video frames. We intend to make our code available to release all the details and support further research.
Appendix B Frequency-based Encoding and Annealing
NeRF [33] has highlighted that a heuristic application of sinusoidal functions to input coordinates, known as “positional encoding”, enables the coordinate-based MLPs to capture higher frequency content, as denoted by:
$\gamma(p)=\left(\sin(2^{0}\pi p),\cos(2^{0}\pi p),\ldots,\sin(2^{L-1}\pi p),\cos(2^{L-1}\pi p)\right)\text{, }$ (9)
where $p$ refers to the point coordinates.
We found that this property also applies to our MLPs that use coordinates of control points as input. This allows the MLPs in the base and motion field to more effectively represent high-frequency information, corresponding to the detailed geometric features. Additionally, when using coordinate-based MLPs to model motion, a significant challenge is capturing both minute and large motions. Following Nerfies [39], we employ a coarse-to-fine strategy that initially targets low-frequency (large-scale) motion and progressively refines the high-frequency (localized) motions. Specifically, we use the following formula to apply weights to each frequency band $j$ in the positional encoding of the MLPs within the motion field.
$w_{j}(\alpha)=\frac{1-\cos(\pi\cdot\text{clamp}(\alpha-j,0,1))}{2},$ (10)
Figure 11: Visual comparison with and without frequency encoding and annealing. The geometry and motion quality degrades when removing annealed frequency-based encoding.
Fig. 11 shows the visual ablation comparisons with and without the frequency encoding and annealing. When removing annealed frequency-based encoding, the geometry and motion quality of the generated animation suffer. To be specific, the butterfly animation in Fig. 11 (row 2) exhibits a lack of geometry details, and the bullfighter animation in Fig. 11 (row 4) shows unreasonable motion, leading to the degradation of the animation quality.
Appendix C Effect Analysis of Control Points
By adjusting the number of control points on the Bézier curves, we can alter the appearance of the generated text animations, thus producing more diverse outcomes. Generally, increasing the number of control points enriches the geometric details, as shown in Fig. 12 (row 2). However, including too many control points may lead to self-intersection of the Bézier curves, causing abrupt and frequent changes of black and white regions, as highlighted by the red circles in Fig. 12 (row 3). Users can choose the sample they desire based on their preferences.
Figure 12: The effect of the number of control points. The first row displays the result generated with the default number of control points extracted by FreeType [8], which is 75. In the second and third rows, the number of control points is increased to 204 and 420 respectively.
Figure 13: Dynamic Typography over different fonts for the same animation sample. The corresponding fonts in the first, second, and third rows are KaushanScript-Regular, Segoe Print, and Roboto-Bold respectively. The animated letter “H” preserves the unique style of each font while faithfully depicting the prompts, allowing it to be seamlessly integrated into the word “HARMONY” under different fonts.
Figure 14: Dynamic Typography over different prompts for the same letter “M” to be animated. Our approach displays different visual effects based on three different prompts, offering a completely different reading experience for the same word.
{CJK*}
UTF8gkai {CJK*}UTF8min
Figure 15: Dynamic Typography over different languages. In the first row, we animate the Chinese character “从” in the word “跟从”, meaning “following” in English. In the second row, we animate the hiragana character 「つ」 from the Japanese word 「つり」, meaning “fishing” in English. This demonstrates the potential of our proposed methodology to generate Dynamic Typography over different languages.
Appendix D Generalizability over Different Fonts, Prompts and Languages
Our method generalizes well in generating Dynamic Typography over different fonts, text prompts, and languages. When the letter to be animated takes different fonts, the generated animation preserves the unique style of each font, as illustrated in Fig. 13. In Fig. 14, by adjusting the input prompts, our method can generate animation with distinct visual effects for the same letter to be animated. Fig. 15 demonstrates two Dynamic Typography samples in Chinese and Japanese, indicating the potential to generalize into different languages.
Appendix E User Study
We conduct a user study to further compare our animation result with three baseline models: DynamiCrafter [57], Gen-2 [46], and LiveSketch [14], using the same dataset as the quantitative analysis. In the study, users are asked to select their preferred sample from four options, each produced by one of the four models. Their selections are guided by the three crucial requirements for Dynamic Typography as described in the method section in the main body, i.e., legibility preservation, semantic alignment, and temporal consistency. Users are required to answer 8 questions that are randomly sampled from the dataset. We collect 62 responses in total.
Figure 16: User Study results. Users are required to select the best animation based on legibility preservation, semantic alignment, and temporal consistency. The pie chart shows the percentage of each method being selected as the best animation.
According to the final result shown in Fig. 16, our animation demonstrates transcendent distinction among the baseline models. Out of $44.76\%$ responses, the animation generated by our method is selected to be the best, indicating that our generated animations best meet the crucial requirements.
Appendix F Failure Case
We observe that, in some samples, the semantic meaning and corresponding motion specified by the user-provided text prompt significantly diverged from the original shape of the letter. In such cases, the model struggles to simultaneously maintain the shape of the letter and convey the vivid semantic information of the text prompt. As a result, the letter either undergoes minimal change in shape, retaining its original form, or it completely loses its original shape, compromising legibility.
Figure 17: Failure case illustration. The first row is generated with the default weight for legibility and structure preservation loss, which suffers from minimal semantic deformation. In the second row, we reduce the weight of these regularization losses, which compromises the legibility.
For example, in Fig. 17, when we incorporate the legibility regularization, the shape of the letter “R” remains unchanged, maintaining its original form while performing the “launch” action. Conversely, when we reduce the weight of the legibility regularization, it transforms completely into the shape of a rocket, losing the characteristic contours of the letter “R”, thus sacrificing the legibility.
Appendix G GPT-4V as Dynamic Typography Designer
As illustrated in Fig. 17, if the user-specified text prompt deviates too much from the chosen letter’s shape, it can hinder the creation of vivid animations. We can utilize the powerful visual and semantic understanding capabilities of Vision Language Models (VLMs) to assist users in selecting appropriate letters and prompts.
In the experiment, we provide current state-of-the-art VLM, GPT-4V [37], with a snapshot of an animation generated, along with the corresponding chosen word, letter, and text prompt as an example to facilitate in-context learning by GPT-4V. Subsequently, we request GPT-4V to design text animations by following the paradigm exemplified in the previous experiment. We require GPT-4V to generate outputs including the word, the selected letter, and the text prompt, explicitly demanding that it considers the similarity between the letter’s original shape and its shape after deformation.
Figure 18: Some results generated by our method based on pairs of prompts and letters that are designed with the assistance of GPT-4V.
We list some samples designed by GPT-4V as follows:
Word: CAT
Chosen Letter: C
Text Prompt: “A cat curls up to sleep.”
Animation Idea: The “C” naturally curls tighter into a circular shape, resembling a cat curling up.
Word: KICK
Chosen Letter: K
Text Prompt: “A soccer player kicks a ball.”
Animation Idea: The angled legs of the “K” mimic the motion of kicking, with one leg drawing back and then striking forward.
Word: SNAKE
Chosen Letter: S
Text Prompt: “A snake slithers through grass.”
Animation Idea: The natural curve of “S” undulates slightly, resembling the slithering movement of a snake.
The generated animations designed by GPT-4V are shown in Fig. 18. We found that GPT-4V has the potential to design proper pairs of prompts and letters that carefully consider the natural shapes of the letters and how they can effectively transform into the desired actions or characteristics with minimal deviation, ensuring the animations are feasible and visually coherent.